OpenStudy (anonymous):

3x + 2y + 5x = 12 so that y is a function of x

OpenStudy (inkyvoyd):

if you want to write an expression such that one variable is a function of the other, you just isolate that variable. So in this case you're trying to solve for y by moving x and constants to the other side. I would begin by combining like terms.

OpenStudy (anonymous):

Ok. It has been many years for me and I am trying to help my grandson. I have forgotten a lot of these terms. if you can break it down for me, I will understand better.

OpenStudy (inkyvoyd):

Okay... so we're trying to solve for y and we expect to see something like y=(some expression with x and constants). I'll refer to y as the dependent and x as the independent variable here, as by definition y is a function of x. so as usual we combine like terms first 2y+8x=12 then we isolate our dependent variable 2y=12-8x then we divide both sides by 2 to get our answer y=6-4x